If cot theta =3,0 in quad III, find sin theta and cos theta
Answer by Alan3354(69443)

If cot theta =3,0 in quad III, find sin theta and cos theta
---------
tan = 1/cot = 1/3
--
tan = y/x --> y = -1, x = -3
r = sqrt(x^2 + y^2) = sqrt(10)
------
sin = y/r = -1/sqrt(10) = -sqrt(10)/10
cos = x/r = -3sqrt(10)/10